OpenStudy (anonymous):

can someone please help me?? Identify the equation of a circle that has a diameter with the end points (6,-8) (10,4)

OpenStudy (anonymous):

A- (x-8)^2 + (y+2)^2 =40 B- (x+8)^2 + (y+2)^2 = 40 c- (x-8)^2 + (y-2)^2 = 40

OpenStudy (anonymous):

Since the circle's diameter is defined by the segment between the points (6,-8) and (10,4) , then it's center is on the center of this segment. The X-coordinate of the circle's center is in the middle between 6 and 10 = 8. The y-coordinate of the circle's center is in the middle between -8 and 4 = -2 So the equation of the circle is: (x - 8)^2 + (y - (-2)) ^2 = 40 (x-8)^2+(y+2)^2=40 ----> Answer A